OpenStudy (anonymous):

Pre-calculus question! if < A=80°, < B=20°, b=12cm then find a??? Side a=________ Can anyone shed some light??This is my last hw question and I don't really know how I would find angle a D: Im sure this is a brainer, but could anyone help? :O Please and thank you!

OpenStudy (anonymous):

remember the sine rule? \[{\sin A\over a}={\sin B\over b}\] when you have an angle and side opposite to it given to ya, check if you can use the sine rule

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

so, a=\[\frac{ sinB }{ b(\sin(A)) }\] =\[\frac{ \sin(20) }{ 12(\sin(80)) }\] =.029 ??? :O

OpenStudy (anonymous):

your re-arrangement of the terms is wrong. try again with the first step

OpenStudy (anonymous):

hmm... ok, a=\[\frac{ \sin(20) }{ \sin(80) }*(12)??\] =4.17 ?? D:

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

why not go step by step? there is no rush. take it slow man. \[ {\sin(80^\circ)\over a}={\sin(20^\circ)\over 12} \]

OpenStudy (anonymous):

ok, I think this is it lol and yea I'm kind of rushing, and not thinking I'm about to go visit family soon and I want to get as much of this done as possible ^^; a=\[\frac{ b }{ sinB }*(sinA)?\]

OpenStudy (anonymous):

I'm about to leave now. I'll be back in the evening and I will see your response then!! Thanks for all your time!

OpenStudy (anonymous):

yes. that is correct.